# Embodied World Models Emerge from Navigational Task in Open-Ended Environments

### v1

@@ Abstract
Understanding how artificial systems can develop spatial awareness and reasoning has long been a challenge in AI research. Traditional models often rely on passive observation, but embodied cognition theory suggests that deeper understanding emerges from active interaction with the environment. This study investigates whether neural networks can autonomously internalize spatial concepts through interaction, focusing on planar navigation tasks. Using Gated Recurrent Units (GRUs) combined with Meta-Reinforcement Learning (Meta-RL), we show that agents can learn to encode spatial properties like direction, distance, and obstacle avoidance. We introduce Hybrid Dynamical Systems (HDS) to model the agent-environment interaction as a closed dynamical system, revealing stable limit cycles that correspond to optimal navigation strategies. Ridge Representation allows us to map navigation paths into a fixed-dimensional behavioral space, enabling comparison with neural states. Canonical Correlation Analysis (CCA) confirms strong alignment between these representations, suggesting that the agent’s neural states actively encode spatial knowledge. Intervention experiments further show that specific neural dimensions are causally linked to navigation performance. This work provides an approach to bridging the gap between action and perception in AI, offering new insights into building adaptive, interpretable models that can generalize across complex environments. The causal validation of neural representations also opens new avenues for understanding and controlling the internal mechanisms of AI systems, pushing the boundaries of how machines learn and reason in dynamic, real-world scenarios.
@@ 1 Introduction
1.1 Background and Motivation
Understanding and constructing world models has long been a central challenge in cognitive science and artificial intelligence [1, 2, 3]. Traditional approaches typically emphasize representing external entities and their interactions, which can support certain levels of prediction and reasoning. However, they often rely on passive perception, in which an agent merely observes the external environment and separates its internal “understanding” from real actions. By contrast, Embodied Cognition theory argues that genuine comprehension arises not only from perception but also from repeated interaction with, and adaptation to, the environment [4, 5]. For instance, concepts such as “up/down” originate from direct bodily experience and gravity, while deeper notions of spatial paths and obstacle avoidance typically take shape through physical action.
@@ 1.1 Background and Motivation
In contemporary AI applications, embodied intelligence is increasingly significant, especially when external annotations are limited and the environment is highly complex [6]. In unsupervised or sparse-reward settings, passive observation often fails to produce the structured spatial understanding necessary for effective decision-making. However, when agents can encode key elements such as direction, distance, and obstacles through repeated environmental interactions, they gain stronger generalization and reasoning capabilities.
Recent developments like Active Inference [1] and Meta-Reinforcement Learning (Meta-RL) [7, 8] have opened new possibilities for embodied cognition. Concurrently, work by Ha & Schmidhuber (2018) has shown how environmental dynamics can be modeled in latent spaces. Building on these theoretical and technical advances, this study explores a fundamental question: Can artificial neural networks autonomously develop sophisticated representations of actions and spatial relationships in embodied tasks such as planar navigation? If the answer is yes, then the resulting “systematic encoding of direction/distance/shape deformation” would extend far beyond basic stimulus-response mappings, with implications for better generalization and interpretability.
1.2 Problem Statement
A key question arises: when an agent influences the environment through its own actions and perceives only local information (as in planar navigation), does it automatically form detailed internal representations of core factors such as direction, obstacles, and path differences? Moreover, do these representations persist as stable, verifiable structures rather than ephemeral memories? If such structures exist, the agent may have truly “internalized” spatial concepts while mastering basic navigation. If not, its learning may remain a shallow policy guided by external rewards, lacking deeper structural knowledge.
The significance of this question is evident. Demonstrating that an agent acquires systematic spatial representations (e.g., direction, distance) through embodied interaction would provide a foundation for higher-level reasoning and explainable decision-making. Rather than simply “remembering to turn left” in response to a specific stimulus, the network would recognize distinctions and connections among spatial concepts such as “up/down.” This broader understanding would also enhance an agent’s ability to transfer to novel environments or adapt to local perturbations, thereby offering valuable benefits in robot navigation, virtual simulation, and biologically inspired intelligent systems.
1.3 Main Contributions
To address these issues, we present three key contributions:
  1. 1.
    Verifying whether neural networks can develop embodied cognitive models of behavior
    We combine Gated Recurrent Units (GRUs) with Meta-Reinforcement Learning (Meta-RL) in a planar navigation task to investigate whether agents can internalize directions, distances, and more complex spatial dimensions, going beyond simple stimulus-response patterns.
  2. 2.
    Introducing Hybrid Dynamical Systems (HDS) analysis and cyclic stimulation
    We unify physical environment states $(Q)$ and neural network hidden states $(X)$ into a single dynamical system $(Q,X)$, enabling us to observe how agent-environment coupling internalizes optimal strategies at a deep level. We also develop a cyclic stimulation method that helps detect and induce system limit cycles—an approach combined with Lyapunov analysis to address the “causal paradox” typical of explainable AI in embodied contexts.
  3. 3.
    Providing multi-perspective evidence: Ridge Representation, CCA, and intervention experiments
    We employ Ridge Representation to embed navigation paths into a fixed-dimensional metric space. We then apply Canonical Correlation Analysis (CCA) to quantify high-dimensional correspondences between neural states and behavioral descriptors, and conduct intervention/control experiments to confirm that altering the most correlated neural dimensions degrades navigation performance, thus establishing a causal link rather than a mere correlation.
Overall, our study theoretically examines how action-perception interactions shape deep conceptual understanding, as posited by embodied cognition, and methodologically shows how hybrid dynamical systems and cyclic stimulation can overcome interpretability challenges posed by agent-environment feedback loops. Empirically, we demonstrate robust encodings of spatial factors such as direction and distance, validated through Ridge Representation, CCA, and targeted interventions. Subsequent sections detail the experimental design, core findings, and broader implications.

### v2

@@ Abstract
Spatial reasoning in partially observable environments has often been approached through passive predictive models, yet theories of embodied cognition suggest that genuinely useful representations arise only when perception is tightly coupled to action. Here we ask whether a recurrent agent, trained solely by sparse rewards to solve procedurally generated planar mazes, can autonomously internalize metric concepts such as direction, distance and obstacle layout. After training, the agent consistently produces near-optimal paths in unseen mazes, behavior that hints at an underlying spatial model.
To probe this possibility, we cast the closed agent–environment loop as a hybrid dynamical system, identify stable limit cycles in its state space, and characterize behavior with a Ridge Representation that embeds whole trajectories into a common metric space. Canonical correlation analysis exposes a robust linear alignment between neural and behavioral manifolds, while targeted perturbations of the most informative neural dimensions sharply degrade navigation performance.
Taken together, these dynamical, representational, and causal signatures show that sustained sensorimotor interaction is sufficient for the spontaneous emergence of compact, embodied world models, providing a principled path toward interpretable and transferable navigation policies.
@@ 1 Introduction
Intelligent behaviour hinges on internal world model—a representation that predicts how the environment will change in response to the agent’s own actions. Early cognitive–computational frameworks attempted to build such models from passive observation, separating perception from control [1, 2, 3]. By contrast, the theory of embodied cognition argues that durable abstractions arise only when perception and action are inseparable: spatial notions such as “up” and “down” do not emerge from merely viewing elevator arrows, but from repeated bodily encounters with gravity and the effort of moving against it [4, 5]. Although modern reinforcement-learning systems can solve intricate navigation tasks, it remains contentious whether their hidden states encode metric structure or merely cache brittle stimulus–response associations.
@@
A key question arises: when an agent actively influences its surroundings while perceiving only a local view, does it automatically develop detailed, stable representations of direction, obstacles and path differences? An affirmative answer would imply genuine internalisation of spatial knowledge; a negative one would suggest that competent performance can still rest on superficial heuristics that may fail outside the training distribution.
We address this question by training a gated-recurrent agent with a meta-reinforcement-learning outer loop [7, 8] on thousands of procedurally generated $10\times 10$ mazes released by an open-ended curriculum generator [6]. Sparse rewards and a $3\times 3$ field of view force the network to accumulate evidence over time and to reuse its hidden state after each goal reset, mirroring the incremental learning posited by embodied-cognition theories. Across 10 000 unseen layouts the agent achieves a $96\%$ reach rate and a final-trial path optimality of $0.92$, indicating that rote memorisation alone cannot explain its behaviour.
To uncover what the network has learned, we (i) model the closed agent–environment loop as a Hybrid Dynamical System, locating limit cycles whose negative Lyapunov exponents mark attractor-level strategies; (ii) embed entire trajectories with a Ridge Representation that places behaviour in a metric space; and (iii) apply Canonical Correlation Analysis to align this space with high-dimensional hidden states. The first ten canonical modes each exceed $0.8$ correlation, and selectively perturbing those neural dimensions doubles path length and halves success, demonstrating causal necessity. These dynamical, representational and causal results show that sustained sensorimotor interaction alone can drive a recurrent network to construct compact, interpretable embodied world models, offering both mechanistic insight into spatial concept formation and a practical diagnostic toolkit for future navigation agents.